# A Comprehensive Study on A Tapered Paul Trap: From Design to Potential Applications

### v1

@@ 1 Introduction
This article begins with an introduction to the design and construction of this tapered ion trap. We then proceed to a numerical field simulation and a 3D trajectory simulation. These simulations allow us to evaluate the performance of our trap. Following this, we provide a detailed description of the experimental setup and present characterizations of this system. Our measurements characterize the axial position dependent radial confinement as a special feature of this system. We also present sideband spectroscopy performed on the $4^{2}S_{1/2}\leftrightarrow 3^{2}D_{5/2}$ quadrupole transition using the electron shelving technique. Finally, we discuss the potential applications of this system in various fields, including quantum thermodynamics, quantum metrology and sensing, and quantum information.
@@ 2 Trap design and field simulation
Figure 2: Pseudopotential simulations: (a) Simplified trap geometry for numerical simulations. The RF blades (blue and red) are driven with a voltage of opposite sign. The endcap electrodes (green) provide confinement along the $z$-direction by applying DC voltages. The compensation electrodes (grey) are responsible for micromotion compensation. (b) The pseudopotential in the $x$-$z$-plane. The black solid and cyan dashed lines illustrate a fitted harmonic potential, showing the axial position dependent confinement.
@@ 3 Experimental setup
Figure 3: Experimental setup of the tapered Paul trap. (a) A helical resonator provides symmetric radio frequency drive to the RF blades avoiding excessive micromotion in the axial direction. DC-voltages $V_{\textrm{D1}}$ and $V_{\textrm{D2}}$ are applied to end-caps. The voltages appled to the compensation electrodes ($V_{\textrm{C1}}$ to $V_{\textrm{C4}}$) are used for micromotion compensation and radial principal axes adjustment. An intensified CCD camera is used to image single ions with a viewing direction perpendicular to the $z$ axis and at 45 degrees with respect to the $x$ and $y$ axis. (b) Top view on the trap: Calcium atoms are provided by a thermal source. The laser system consists of ionization lasers (at 422 nm and 375 nm) with cooling (at 397 nm) and repumping lasers (at 854 nm and 866 nm) for Doppler cooling. A quantization axis is defined by applying 3 G magnetic field by a pair of permanent magnet rings counter-propagating with a narrow band 729 nm laser used for the $4^{2}S_{1/2}$ to $3^{2}D_{5/2}$ quadrupole transition. (c) Relevant energy levels and optical transitions of the 40Ca+ ion.
Figure 3 (a) shows the electric connections to the Paul trap. The dynamic confinement in the radial directions ($x$, $y$) is provided by RF voltages ($+\widetilde{V}_{\mathrm{RF}}$, $-\widetilde{V}_{\mathrm{RF}}$) applied to the two opposite-positioned pairs of blades, respectively. An RF signal at 11.17 MHz with a power of -17 dBm is amplified by a 5-watt power amplifier with a gain of 46 dB. The amplified signal is then sent to a helical resonator for further amplification and impedance matching to the trap. This helical resonator is specifically designed with a $\lambda/2$ geometry [27], whose middle point is grounded with two open ends connected to the two pairs of RF blades. This configuration provides two RF voltages with a $\pi$ phase difference that is necessary to avoid excessive axial micromotion. The RF voltages are probed by voltage dividers composed of two capacitors for ease of characterization and adjustment. Through careful adjustment of the sliding contact on the coil and the tunable capacitors on the helical resonator, we tune to an impedance-matched frequency of 11.17 MHz. The phase difference of the two RF signals is measured to be 179.51(±$\pm 0.28$) degrees with a relative amplitude difference of 1.29%. The axial confinement is achieved by applying DC voltages ($V_{\mathrm{D1}}$ and $V_{\mathrm{D2}}$) to the endcap electrodes. Additionally, four auxiliary electrodes are implemented around the Paul trap (see Fig. 1 in yellow). These electrodes are supplied with four independent DC voltage sources, which allow for rotating radial principal axes and compensation of the micromotion [28].
@@ 6 Discussion and Outlook
With this paper, we want to draw attention to our tapered Paul-trap platform that allows utilizing its special features, such as mechanical nonlinearities, to implement novel schemes for quantum information, to advance metrology and sensing methods, and to push forward quantum thermal machines for probing thermodynamics in the quantum regime. A detailed description of the setup presented in this article allows analytical and numerical investigations of the trap and provides a sound basis for designing, optimizing, and tailoring properties of the system for new applications and basic research.
Motional modes of trapped ions are used as a quantum bus for quantum information processing [34]. In the tapered trap, radial modes for neighboring ions have different frequencies, allowing local addressing in frequency domain. This enables new ways of implementing quantum simulation and gate operations by selectively addressing the modes in frequency space, which allows for global laser irradiation of an ion crystal.
The mechanical nonlinearity opens up new possibilities to advance quantum sensing and metrology through a range of effects, such as multistability, stochastic and vibrational resonances [35]. Similar to demonstrations with nanomechanical oscillators  [36], where a broadband noise has led to significant signal amplification, our system has shown similar behavior [17] through vibrational resonance. Additionally, combining quantum resources such as squeezing with nonlinearities could further increase the sensitivity similar to effects predicted with quantum thermal machines based on trapped ions [37].
Thermodynamics in the quantum regime is a new active research field [38, 39]. Trapping ions in tapered traps have shown pioneering results with the first single ion heat engine being realized [18] and parameters for work extraction at the Curzon-Ahlborn efficiency limit being predicted [21]. Quantum resources such as squeezing are expected to increase the efficiency of heat extraction [37] and transient non-confining potentials can be used for speeding up a single ion heat pump through the use of shortcuts to adiabaticity [40]. The single-ion heat engine can also be used as a sensitive thermal probe [41] and the sensitivity can be enhanced through the application of squeezing. The tapered trap design also potentially allows for implementing local non-classical baths combining dissipative state preparation [42] with frequency space addressing of the local radial modes, which is of fundamental interest in quantum resource theory [43].

### v2

@@ 1 Introduction
This article begins with an introduction to the design and construction of this tapered ion trap. We then proceed to a numerical field simulation and a 3D trajectory simulation. These simulations allow us to evaluate the performance of our trap. Following this, we provide a detailed description of the experimental setup and present characterizations of this system. Our measurements characterize the axial position dependent radial confinement as a special feature of this system. We also present sideband spectroscopy performed on the $4^{2}S_{1/2}\leftrightarrow 3^{2}D_{5/2}$ quadrupole transition of 40Ca+ ions using the electron shelving technique. Finally, we discuss the potential applications of this system in various fields, including quantum thermodynamics, quantum metrology and sensing, and quantum information.
@@ 2 Trap design and field simulation
Figure 2: Pseudopotential simulations: (a) Simplified trap geometry for numerical simulations. The RF blades (blue and red) with an axial length of 4 mm are driven with a voltage of opposite sign. The endcap electrodes (green) spaced by 4.8 mm provide confinement along the $z$-direction by applying DC voltages. The compensation electrodes (grey) are responsible for micromotion compensation (diagonal distance 17 mm and 2 mm diameter). (b) The pseudopotential in the $x$-$z$-plane. The black solid and cyan dashed lines illustrate a fitted harmonic potential, showing the axial position dependent confinement.
@@ 3 Experimental setup
Figure 3: Experimental setup of the tapered Paul trap. (a) A helical resonator provides a symmetric radio frequency drive to the RF blades avoiding excessive micromotion in the axial direction. DC-voltages $V_{\textrm{D1}}$ and $V_{\textrm{D2}}$ are applied to end-caps. The voltages applied to the compensation electrodes ($V_{\textrm{C1}}$ to $V_{\textrm{C4}}$) are used for micromotion compensation and radial principal axes adjustment. An intensified CCD camera is used to image single ions with a viewing direction perpendicular to the $z$ axis and at 45 degrees with respect to the $x$ and $y$ axis. (b) Top view on the trap: Calcium atoms are provided by a thermal source. The laser system consists of ionization lasers (at 422 nm and 375 nm) with cooling (at 397 nm) and repumping lasers (at 854 nm and 866 nm) for Doppler cooling. A quantization axis is defined by applying 3 G magnetic field by a pair of permanent magnet rings counter-propagating with a narrow band 729 nm laser used for the $4^{2}S_{1/2}$ to $3^{2}D_{5/2}$ quadrupole transition. (c) Relevant energy levels and optical transitions of the 40Ca+ ion.
Figure 3 (a) shows the electric connections to the Paul trap. The dynamic confinement in the radial directions ($x$, $y$) is provided by RF voltages ($+\widetilde{V}_{\mathrm{RF}}$, $-\widetilde{V}_{\mathrm{RF}}$) applied to the two opposite-positioned pairs of blades, respectively. An RF signal at 11.17 MHz from an analog signal generator is amplified by a 5-watt power amplifier with a gain of 46 dB. The amplified signal is then sent to a helical resonator for further amplification and impedance matching to the trap. It is specifically designed with a $\lambda/2$ geometry [27], whose middle point is grounded with two open ends connected to the two pairs of RF blades. This configuration provides two RF voltages with a $\pi$ phase difference that is necessary to avoid excessive axial micromotion. The RF voltages are probed by voltage dividers composed of two capacitors with a ratio of 118 for ease of characterization and adjustment. Through careful adjustment of the sliding contact on the coil and the tunable capacitors on the helical resonator, we tune to an impedance-matched frequency of 11.17 MHz, where the quality factor of this helical resonator together with the Paul trap is measured by a vector network analyzer to be about 46. Through analyzing the probed RF signal on the oscilloscope, the phase difference of the two RF signals is measured to be 179.51(±$\pm 0.28$) degrees with a relative amplitude difference of 1.29%. The amplitudes of both probed RF signals are around 0.8 V. The RF voltages on the blade-shaped RF electrodes have amplitudes of around 95 V. The axial confinement is achieved by applying DC voltages ($V_{\mathrm{D1}}$ and $V_{\mathrm{D2}}$) to the endcap electrodes. Additionally, four auxiliary electrodes are implemented around the Paul trap (see Fig. 1 in yellow). These electrodes are supplied with four independent DC voltage sources, which allow for rotating radial principal axes and compensation of the micromotion [28].
@@ 6 Discussion and Outlook
A detailed description of the setup presented in this article allows analytical and numerical investigations of the trap and provides a sound basis for designing, optimizing, and tailoring properties of the system for new applications and basic research. We want to draw attention to our tapered Paul trap platform and spark interest to utilize its special features, such as mechanical nonlinearities, to implement novel schemes for quantum information, to advance metrology and sensing methods, and to push forward quantum thermal machines for probing thermodynamics in the quantum regime.
As an example, in the tapered trap, radial modes for neighboring ions have different frequencies, allowing local addressing in the frequency domain. This enables new ways of implementing quantum simulation and gate operations, such as selectively addressable single-qubit operations in frequency space and exploiting the radial-axial coupling for performing selective two-qubit operations, extending previously utilized quantum bus methods [34, 35].
The mechanical nonlinearity opens up new possibilities to advance quantum sensing and metrology through a range of effects, such as bistability, stochastic and vibrational resonances [36]. For instance, coherent signal amplification was achieved by using bistable nanomechanical oscillators [37], where the use of noise has led to significant signal amplification. Similarly, we amplified a signal at zeptonewton-level by exploiting vibrational resonances [17]. Additionally, combining quantum resources such as squeezing with nonlinearities could further increase the sensitivity, similar to effects predicted with quantum thermal machines based on trapped ions [38].
Furthermore, thermodynamics in the quantum regime is a new active research field [39, 40], where tapered Paul traps have shown pioneering results with the first single ion heat engine being realized [18] and parameters for work extraction at the Curzon-Ahlborn efficiency limit have been predicted [21]. Quantum resources such as squeezing are expected to increase the efficiency of heat extraction [38] and transient non-confining potentials can be used for speeding up a single ion heat pump through the use of shortcuts to adiabaticity [41]. The single-ion heat engine can also be used as a sensitive thermal probe and the sensitivity can be enhanced through the application of squeezing [42]. The tapered trap design also potentially allows for implementing local non-classical baths combining dissipative state preparation [43] with frequency space addressing of the local radial modes, which is of fundamental interest in quantum resource theory [44].